OpenStudy (anonymous):

open parentheses 8 plus square root of 5 close parentheses open parentheses 8 minus square root of 5 close parentheses

OpenStudy (anonymous):

\[(8+\sqrt{5})(8-\sqrt{5})\]

OpenStudy (anonymous):

Is That your question?

OpenStudy (jdoe0001):

\(\bf (a-b)(a+b) = a^2-b^2\)

OpenStudy (anonymous):

Follow the formula provided by @jdoe0001 and you'll get your answer.

OpenStudy (anonymous):

59?